A certain weak base has a Kb of 9.00 x 10-7. What concentration of this base will produce a pH of 10.06? concentration:
Answer #1

Let the concentration of B c molar

use:

pH = -log [H+]

10.06 = -log [H+]

[H+] = 8.71*10^-11 M

use:

[OH-] = Kw/[H+]

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

[OH-] = (1.0*10^-14)/[H+]

[OH-] = (1.0*10^-14)/(8.71*10^-11)

[OH-] = 1.148*10^-4 M

B dissociates as:

B +H2O -----> BH+ + OH-

c 0 0

c-x x x

Kb = [BH+][OH-]/[B]

Kb = x*x/(c-x)

9*10^-7 = 1.148*10^-4*1.148*10^-4/(c-1.148*10^-4)

c-1.148*10^-4 = 1.465*10^-2

c=1.48*10^-2 M

Answer: 0.0148 M